OpenStudy (anonymous):

Solve by factoring: 6x²-3x-30=0 please help I get to a part where i cant factor evenly. I will give a medal.

OpenStudy (anonymous):

The most sure way to factor an equation like this is to use the quadratic equation: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

Isn't that not factoring?

OpenStudy (anonymous):

It is of the format such that: \[ax^2 + bx + c\] In your problem we have a=6 b=-3 and c= -30

OpenStudy (anonymous):

it will give you something of x= -a and x = +b the factors therefore are (x+a)(x-b) note the sign of 'a' and 'b' change

OpenStudy (anonymous):

so i should use the quadratic formula?

OpenStudy (anonymous):

I used bad variables for -a and +b,,, call them x = -u and x = +v so we have (x+u)(x-v)

OpenStudy (anonymous):

yes, however this one DOES factor nicely, although it may be hard to see at first

OpenStudy (anonymous):

(3x + 6 ) (2x -5)

OpenStudy (anonymous):

This doesn't work i got a decimal for one of my answers for the quadratic formula

OpenStudy (anonymous):

im assuming your answers are x = -2 and x = +2.5

OpenStudy (anonymous):

the quadratic equation gives you values of 'x' which make the equation = 0. when you factor an equation, you get (x+c)(x+d)=0 so when x = -c or x = -d it all equals 0.

OpenStudy (anonymous):

The quadratic equation jumps all the way to the end: x = -c or x = -d so to get the factors you simply make it (x+c) = 0 or (x+d) = 0 thus the factors are (x+c)(x+d)=